Assume the Knight Corporation is considering the acquisition of Day Inc. The expected earnings per share for the Knight Corporation will be $6 with or without the merger. However, the standard deviation of the earnings will go from $1.80 to $1.32 with the merger because the two firms are negatively correlated.


a. Compute the coefficient of variation for the Knight Corporation before and after the merger. (Do not round intermediate calculations and round your answers to 2 decimal places.)
  


b. Comment on the possible impact on Knight’s postmerger P/E ratio, assuming investors are risk-averse.
  

Risk-averse investors are being offered    risk and may assign a P/E ratio to postmerger earnings.
Answer #1

1.

Before merger=1.80/6=0.30

After merger=1.32/6=0.22

2.

Risk averse investors are being offered low risk and may assign a higher PE ratio to postmerger earnings